Solve the system of equations by elimination.

2x+12y=-50
10x+3y=35
please help

Answer by mananth(16946)

2x+12y=-50......................1
10x+3y=35.......................2
...
multiply equation 1 by -5
-10x-60y=250
add it to equation 2
10x-10x+3y-60y=250+35
-57y=285
/-57
y=-5
plug value of y in equation 1
2x-60=-50
2x=10
x= 5
...
m.ananth@hotmail.ca